### v1

# End-To-End Causal Effect Estimation from Unstructured Natural Language Data

Abstract
Knowing the effect of an intervention is critical for human decision-making, but current approaches for causal effect estimation rely on manual data collection and structuring, regardless of the causal assumptions. This increases both the cost and time-to-completion for studies. We show how large, diverse observational text data can be mined with large language models (LLMs) to produce inexpensive causal effect estimates under appropriate causal assumptions. We introduce NATURAL, a novel family of causal effect estimators built with LLMs that operate over datasets of unstructured text. Our estimators use LLM conditional distributions (over variables of interest, given the text data) to assist in the computation of classical estimators of causal effect. We overcome a number of technical challenges to realize this idea, such as automating data curation and using LLMs to impute missing information. We prepare six (two synthetic and four real) observational datasets, paired with corresponding ground truth in the form of randomized trials, which we used to systematically evaluate each step of our pipeline. NATURAL estimators demonstrate remarkable performance, yielding causal effect estimates that fall within 3 percentage points of their ground truth counterparts, including on real-world Phase 3/4 clinical trials. Our results suggest that unstructured text data is a rich source of causal effect information, and NATURAL is a first step towards an automated pipeline to tap this resource.
1 Introduction
Estimating the causal effects of interventions is time consuming and costly, but the resulting outcomes are precious. Health agencies around the world often require randomized controlled trial (RCT) data to approve medical interventions. Clinical trials are key contributors to large R&D costs for drug developers [27]. Natural experiments are another source of rich interventional data, but they may not always exist or have enough data relevant to a given causal hypothesis  [11].
When treatment randomization is infeasible, observational data can be used to identify average treatment effects (ATEs) [45], under common assumptions, e.g., no unobserved confounding. Such data is abundant but even when the necessary assumptions are satisfied, it must be structured (i.e., the outcomes, treatments, and relevant covariates must be defined, recorded, and tabulated) before it becomes amenable to computational analyses.
Yet, unstructured observational data presents unique opportunities for cheaper, more accessible, and potentially even better [29] effect estimation. For example, thousands of people living with diabetes choose to share their experiences with treatments on online patient forums. Some of their posts contain rich descriptions of daily lives, the drugs they have been prescribed, the treatment responses and side effects, as well as pre-treatment information like age and sex. Their posts contain their lived experiences including evidence of an outcome in an observational experiment, albeit in an unstructured form. Other potential sources of rich unstructured, observational data include newspaper classifieds, police reports, social media, and clinical reports. Despite being collected for a myriad of purposes, researchers have often turned to such data to test hypotheses since: (i) unstructured data does not require restrictive data collection designs, e.g., measurement choice, and can admit many different post-hoc analyses; (ii) the reported outcomes may reflect what matters to subjects better than standard outcome measures; (iii) value may be recouped from outcomes that would otherwise be lost; (iv) there may be more unstructured data available on underserved or marginalized populations. Figure 1 contrasts our setting with previous works using randomized or structured observational data.
This work asks a simple question: How can we use large language models to automate treatment effect estimation using freely available text data? We introduce NATURAL, a family of text-conditioned estimators that addresses this by performing NATural language analysis to Understand ReAL effects.
At a high level, the steps required to compute NATURAL estimators are as follows. Given an observational study design and a dataset of natural language reports, filter for reports that are likely to conform to the experimental design. Then, using a large language model (LLM), extract the conditional distribution of structured variables of interest (outcome, treatment, covariates) given the report. Finally, use the conditionals to compute estimators of the ATE, using classical strategies such as inverse propensity score weighting and outcome imputation.
NATURAL is a data-driven pipeline. It leverages and relies on the LLM in a manner that mimics the learning task it was trained for: providing parametric approximations to conditional distributions. As in all observational studies, the validity of NATURAL also depends on prior causal knowledge about the task. Expert knowledge is required to define appropriate covariates and confirm that they satisfy the necessary assumptions for effect estimation. However, we anticipate that NATURAL estimators could be developed under other structural assumptions (e.g. instrumental variables) as well.
Figure 1: When compared to experimental and other observational studies, NATURAL has lower costs and provides greater diversity in cohort selection, for causal effect estimation.
The core contributions of our work are:
  • We derived NATURAL ATE estimators based on classical estimators of the ATE, like inverse propensity score weighting and outcome imputation. NATURAL estimators operate on entirely unstructured data under two novel data-access assumptions.
  • We implemented NATURAL estimators using an LLM-based pipeline.
  • We developed six observational datasets to systematically evaluate parts of this pipeline: two synthetic datasets constructed using marketing data, and four clinical datasets curated from public (pre-December 2022) migraine and diabetes subreddits from the Pushshift collection [6].
  • For each dataset, we treated the ATE from a corresponding real-world completely randomized experiment (CRE) as ground truth. Remarkably, our predicted ATEs all fell within 3 percentage points of the ground truth ATEs, a potential cost savings of many millions of dollars.
1.1 Related Work
Leveraging natural language data [39] to support causal claims is pervasive in applied research [40, 12]. Our work falls under the broad umbrella of accelerating the identification of real-world evidence (RWE) [37]. For instance, in the context of healthcare, RWE supports not only drug repurposing, but also post-market safety evaluations — its most common application. NATURAL expands the boundaries of how quickly one can obtain and validate such real-world evidence from observational data [38].
The use of natural language data in causal inference comes in different flavors: i) using text to measure confounders [21], ii) using text to measure causal effect outcomes [13], or iii) producing interpretable causal features from text [13, 5], e.g., what words are more likely to explain the cause of an event. NATURAL distinguishes itself from these lines of research in two ways: i) NATURAL does not require any curated task-specific training data (it is zero-shot), and ii) NATURAL is not interested in how the text itself, i.e., its words, relate to the causal problem —that is, we are only leveraging the model’s ability to predict the distribution of a specified variable conditional on the input text. We highlight that our work lies distinct from research at the intersection of text and causality that has studied the ability of language models to infer latent variables (that are implied but not explicitly identified in text data) [33, 12]. Rather, we require the precise specification of covariates to condition on – we view this as being crucial to creating a more direct way for an end user to verify the validity of information extracted with our approach.
Prior works have also leveraged LLMs in a black-box fashion for causal tasks by querying the model for causal statements. In the context of causal discovery, users directly ask for the existence of cause-and-effect relationships, e.g.,“Does changing the age of an abalone causes a change in its length?” [23, 30, 3, 4, 41, 19, 5]. Due to the large amount of training data, it is possible that the model learns to apply a causal model described in the training data and answer causal questions with it [32, 44]. The issue with this approach is i) the user is limited to the causal models observed in training, ii) the user is not aware of which causal model they are using, and iii) the queries tend to present high prompt sensitivity [26]. Finally, we note that a recent work created a benchmark and showed how LLMs struggle to distinguish pairwise correlation from causation [18], while another shows that checking causal relationships in a pairwise manner can lead to invalid causal graphs [42].
2 Preliminaries
We are interested in estimating the causal effect of a treatment relative to either another treatment or no treatment in a population of interest. More precisely, we consider treatments $t\in\{0,1\}$ and the corresponding potential outcomes $Y(1)$ and $Y(0)$ under each treatment. We wish to compute the quantity $\tau:=\mathbb{E}[Y(1)-Y(0)]$, often referred to as Average Treatment Effect (ATE). Sometimes, $Y(0)$ may correspond to no treatment (control). Throughout this work, we assume binary treatments and outcomes in the Neyman-Rubin causal model. We provide a full list of notation in appendix A.
A Completely Randomized Experiment (CRE) with $n$ participants requires no prior causal knowledge. In a CRE, the treatment assignment vector $(\tilde{T}_{i})_{i=1}^{n}$ is a random permutation of $n_{1}$ ones and $n-n_{1}$ zeros sampled independently of the outcomes. In this case, the difference-in-means $\tfrac{1}{n_{1}}\sum_{i=1}^{n}\tilde{T}_{i}Y_{i}(1)-\tfrac{1}{n-n_{1}}\sum_{i=1}^{n}(1-\tilde{T}_{i})Y_{i}(0)$ provides us with an unbiased estimate of $\tau$.
Despite the indisputable necessity of CREs in high-stakes settings, it is often expensive and/or infeasible to have complete control over the treatment assignment. Instead, observational data is more readily available. Observational data often contains spurious correlations between the observed treatment $T$ and the observed outcome $Y=TY(1)+(1-T)Y(0)$ through a common cause (confounder). Typically, this confounding is formalized as a variable $X$, which we assume to be discrete throughout this work, representing covariates associated with each individual. Given i.i.d. samples $\{(X_{i},T_{i},Y_{i})\}_{i=1}^{n}$ from the target population, standard causal inference techniques can correct for confounding bias and provide consistent estimates of $\tau$ under Assumptions 1 and 2:
Assumption 1 (Strong Ignorability.)
The potential outcomes are independent of treatment assignments conditional on covariates, i.e., $(Y(0),Y(1))\perp\!\!\!\perp T|X.$
Assumption 2 (Positivity.)
For every treatment $t$ and covariate set $x$, $10<P(T=t\mid X=x)<$ 1 .
Following are two classical estimators of the ATE $\tau$ from observational data, each of which rely on $X$ satisfying Assumptions 1 and 2. We refer the reader to Ding, [10] for further details.
Inverse Propensity Score Weighting (IPW).
The propensity score is the conditional probability of receiving a treatment given the observed features, i.e., $e(x)=P(T=1|X=x)$. The IPW estimator is given by
$\hat{\tau}_{\text{IPW}}=\frac{1}{n}\sum_{i=1}^{n}\frac{T_{i}Y_{i}}{\hat{e}(X_{i})}-\frac{(1-T_{i})Y_{i}}{1-\hat{e}(X_{i})},$ (1)
where $\hat{e}(x)$ is an approximation of $P(T=1\mid X=x)$. When $\hat{e}(x)$ is the true propensity score, $\hat{\tau}_{\text{IPW}}$ is an unbiased estimator of $\tau$. When $\hat{e}(x)$ is estimated as empirical probability, $\hat{\tau}_{\text{IPW}}$ is consistent.
Outcome Imputation (OI).
Outcome Imputation learns a model to impute outcomes from features and treatment and then marginalizes away the features to estimate $\tau$ with
$\hat{\tau}_{\text{OI}}=\frac{1}{n}\sum_{i=1}^{n}\hat{\tau}(X_{i},1)-\hat{\tau}(X_{i},0),$ (2)
where $\hat{\tau}(x,t)$ approximates $P(Y=1\mid X=x,T=t)$. Note that if $\hat{\tau}(x,t)$ is an unbiased estimation of this quantity, $\hat{\tau}_{\text{OI}}$ is an unbiased estimator of $\tau$.
3 NATURAL estimators of the ATE
Both CRE and observational studies require direct access to tabulated data $(X_{i}$, $T_{i}$, $Y_{i})$ for every individual $i$. Our NATURAL estimators on the other hand estimate the ATE from observational, unstructured natural language data in the form of language reports $R_{i}$. In addition to Assumptions 1 and 2, NATURAL estimators require the following assumptions to guarantee their consistency.
Assumption 3 (Natural language report data.)
The target population is described by an observational data-generating process $P(X,T,Y,R)$ of data $(X,T,Y)$, which satisfies Assumptions 1 and 2 and is jointly distributed with a random natural language string $R$, called a report. We assume access to an i.i.d. sample of reports $\{R_{i}\}_{i=1}^{n}$ from the marginal of this process.
Assumption 4 (Access to the true observational conditional over $(X,T,Y)$.)
We can either (i) compute the conditional $P(X=x,T=t,Y=y|R=r)$ of the true data-generating process, or (ii) we can sample from $P(X=x|R=r)$ and compute $P(T=t,Y=y|R=r,X=x)$.
Intuitively, these assumptions give NATURAL indirect access to $(X,T,Y)$ through $R$. They can be weak or strong, depending on the definition of the reports $R$. On the one hand, if reports are copies of the observational data, i.e., $R=(X,T,Y)$, then Assumption 4 is trivial to satisfy. On the other hand, if reports are all the constant, empty string, $R=\epsilon$, then Assumption 4 guarantees that we have full access to the true observational joint density function over $(X,T,Y)$, which is a strong assumption. We consider how we might satisfy these assumptions in practice in the next section. Here, we assume that they hold and develop a series of consistent estimators of the ATE.
NATURAL Full. Given $\{R_{i}\}_{i=1}^{n}$ and $P(X=x,T=t,Y=y|R=r)$, we can construct an idealized version of NATURAL. Let us start by noting that the law of total expectation gives us
$\tau=\mathbb{E}_{X,T,Y}\left[\frac{TY}{e(X)}-\frac{(1-T)Y}{1-e(X)}\right]=\mathbb{E}_{R}\left[\mathbb{E}_{X,T,Y|R}\left[\frac{TY}{e(X)}-\frac{(1-T)Y}{1-e(X)}\right]\right].$ (3)
A Monte Carlo estimate over reports is given by
$\hat{\tau}_{\text{N-Full}}=\frac{1}{n}\sum_{i=1}^{n}\sum_{x,t,y}P(X=x,T=t,Y=y|R_{i})\left[\frac{ty}{\hat{e}_{\text{N-Full}}(x)}-\frac{(1-t)y}{1-\hat{e}_{\text{N-Full}}(x)}\right],$ (4)
which further approximates $\hat{e}_{\text{N-Full}}(x)$ from the given conditional. We used eq. 7 below. We note that $\hat{\tau}_{\text{N-Full}}$ above is derived using the form of IPW, but the same estimator can be derived from OI, as shown in appendix B.
The estimator $\hat{\tau}_{\text{N-Full}}$ above relies on enumerating all possible values of $(X,T,Y)$, making it computationally expensive for high-dimensional $X$. Below, we present two hybrid versions of our method which combine sampling of some variables and computation of conditional probabilities of others.
NATURAL IPW. To construct our hybrid estimator, we augment the data $\{R_{i}\}_{i=1}^{n}$ by sampling from $P(X|R_{i})$ independently for each report $R_{i}$. This gives us a dataset $\{(R_{i},X_{i})\}_{i=1}^{n}$ drawn i.i.d. from $P(X,R)$ by Assumption 4. Then, our hybrid estimator is derived from the form of IPW as follows:
$\tau$ $=\mathbb{E}_{R,X}\left[\mathbb{E}_{T,Y|R,X}\left[\frac{TY}{e(X)}-\frac{(1-T)Y}{1-e(X)}\right]\right],$ (5)
$\hat{\tau}_{\text{N-IPW}}$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{(t,y)\in\mathcal{T}\times\mathcal{Y}}P(T=t,Y=y|R_{i},X_{i})\left[\frac{ty}{\hat{e}_{\text{N-IPW}}(X_{i})}-\frac{(1-t)y}{1-\hat{e}_{\text{N-IPW}}(X_{i})}\right],$ (6)
where $\hat{e}_{\text{N-IPW}}(x)$ is consistently estimated in the following manner:
[…]
NATURAL OI. Similarly inspired by the OI estimator in equation 2, we have for $t\in\{0,1\}$,
$P(Y=1\mid T=t,X=x)=\frac{\mathbb{E}_{R,X,T}\big{[}P(Y=1|R,X,T)\mathbb{I}\!\left(X=x,T=t\right)\big{]}}{\mathbb{E}_{R,X,T}\big{[}\mathbb{I}\!\left(X=x,T=t\right)\big{]}}.$ (8)
Thus, for our hybrid OI estimator, we augment the data $\{R_{i}\}_{i=1}^{n}$ by sampling from $P(X,T|R_{i})$ independently for each report $R_{i}$. This gives us a dataset $\{(R_{i},X_{i},T_{i})\}_{i=1}^{n}$ drawn i.i.d. from $P(R,X,T)$ by Assumption 4. Then, our consistent outcome predictor is given by
$\hat{\tau}_{\text{N-OI}}(x,t)=\frac{\sum_{i=1}^{n}P(Y=1|R_{i},X_{i},T_{i})\mathbb{I}\!\left(X_{i}=x,T_{i}=t\right)}{\sum_{i=1}^{n}\mathbb{I}\!\left(X_{i}=x,T_{i}=t\right)},$ (9)
and the final estimator is given by:
$\hat{\tau}_{\text{N-OI}}=\frac{1}{n}\sum_{i=1}^{n}\hat{\tau}_{\text{N-OI}}(X_{i},1)-\hat{\tau}_{\text{N-OI}}(X_{i},0).$ (10)
NATURAL Monte Carlo. Further in the direction of sampling more variables, we can obtain samples $(X_{i},T_{i},Y_{i})$ from $P(X,T,Y|R_{i})$ and compute a Monte Carlo estimate, $\hat{\tau}_{\text{N-MC}}$. The set of samples $\{(X_{i},T_{i},Y_{i})\}_{i=1}^{n}$ constitute a tabular dataset which can be plugged into a standard ATE estimator like IPW or OI, as described in section 2. We refer to these sample-only estimators as N-MC IPW and N-MC OI, respectively.
4 Implementing NATURAL estimators with Large Language Models
LLMs are trained on vast datasets of real-world data, e.g., [2], which likely contain records of data generated by processes that are consistent with Assumption 3. Because LLMs can learn well-calibrated conditionals [20], our hypothesis is that LLMs can be prompted to approximate the conditionals required by Assumption 4 for real-world causal effect questions of interest. Our LLM implementation of NATURAL estimators is built on this hypothesis to try to satisfy Assumptions 3 and 4 (Assumptions 1 and 2 must be guaranteed by a domain expert). We defer exact prompts to appendix D, a full worked example to appendix C, and a discussion of the limitations of our approach to the next section. Figure 2 summarizes our pipeline.
Figure 2: Our pipeline leverages LLMs to curate data that can be plugged into natural language conditioned estimators for average treatment effects.
Filtering to match Assumption 3. For our real-data clinical settings, our first goal is to produce a dataset of i.i.d. reports $R_{i}$ that are very likely to be jointly distributed with the random variables $(X_{i},T_{i},Y_{i})$ of a specific observational study of interest. Given a study of interest and a dataset of real-world reports that are potentially relevant to the study, we pass it through a sequence of filters with increasing detail and strictness:
  1. (i)
    Initial filter. Inspired by other work with social media data [1, 35], we first use deterministic rules to filter out uninformative reports: posts that were removed, are too short, have "bot" in the author’s name, have no mention of any keyword related to the study, etc.
  2. (ii)
    Filter by relevance. We prompt an LLM to determine whether each report contains information that would make it relevant to the study. We remove reports that are deemed irrelevant.
  3. (iii)
    Filter by treatment-outcome. We ensure that each report pertains specifically to the treatments and outcomes of interest. We do so by prompting an LLM to extract only treatment and outcome information, and retaining only the posts that are deemed to both mention one of the treatments in question and also contain outcome information.
  4. (iv)
    Filter known covariates by inclusion criteria. We are sometimes interested in ATEs over populations defined by constraints on pre-treatment covariates $X_{i}$ known as inclusion criteria. In such cases, we included a filter to enforce inclusion criteria. Managing inclusion criteria is complicated by the fact that many reports $R_{i}$ contain no or partial information about covariates that are required to verify inclusion. So, in this filtering step, our goal was to ensure that the final set of reports have non-zero probability of matching the inclusion criteria. We begin by prompting an LLM to extract the full set of covariates $X_{i}$, following constraints on the possible values each attribute can take, but we allow the LLM to extract Unknown if it is impossible for the LLM to determine the value of a covariate. We then remove reports, if any of the non-Unknown covariates are determined to fail their inclusion criteria. We found the JSON-mode made available for generation by certain LLM APIs, like GPT-4, to suffice for this task; however more involved strategies for constrained generation are also possible [43, 47].
Sampling from and computing conditional probabilities to match Assumption 4. Given a set of reports $\{R_{i}\}_{i=1}^{n}$ that pass the filtering stage above, our next steps use LLMs to extract the samples and conditionals $P_{\text{LLM}}(X,T,Y\mid R)$, required to compute NATURAL estimators. For each $R_{i}$, we:
  1. (v)
    Extract covariates, both known and unknown. We run a final covariate extraction by prompting an LLM to determine the full set of covariates $X_{i}$ from the report $R_{i}$, subject to the constraint that $X_{i}$ satisfies the inclusion criteria. In contrast to (iv), we ask the LLM to guess the values of Unknown covariates. We verified that this second extraction agreed exactly with the first extraction (iv) on the known covariates (i.e., the ones that were not extracted as Unknown in the first extraction). We contrast the empirical distributions of these known and unknown/guessed covariates for our experiments in section F.1.
  2. (vi)
    Infer conditionals. Given $\{R_{i},X_{i}\}_{i=1}^{n}$ from the previous steps, we compute the probabilities $P_{\text{LLM}}(T=t,Y=y|R_{i},X_{i})$ by prompting an LLM that makes log-probabilities accessible, like LLAMA2. Specifically, we ask an LLM to answer questions about $T,Y$ given access to $R_{i},X_{i}$, and we score every possible answer $T=t,Y=y$ using the LLM log-probabilities. We exponentiate and renormalize these scores across the space of possible realizations to obtain a valid probability distribution.
To manage inclusion criteria, our LLM pipeline requires three additional technical conditions, which we discuss in more detail in appendix G: (i) inclusion criteria define a box, i.e., a separate criterion is specified for each covariate dimension, (ii) the event that the set of Unknown in (iv) satisfy their inclusion criteria is conditionally independent of the report and the known covariates given the event that the knowns satisfy their criteria, and (iii) the set of known and unknown covariates is conditionally independent of the treatment and outcome given the report and the covariate values. We used these to avoid an expensive estimation or a data-wasteful rejection sampling step, but these assumptions are not required in principle by NATURAL estimators.
Nevertheless, while our empirical results are remarkably consistent with the correctness of our pipeline, we cannot formally guarantee that it satisfies Assumptions 3 and 4. The final outcome of this pipeline is a dataset $\{R_{i},X_{i}\}_{i=1}^{n}$ and a set of conditionals $P(T=t,Y=y|R_{i},X_{i})$ that can be plugged into the hybrid NATURAL estimators in section 3 to predict ATEs. Therefore, we see this as a first implementation of NATURAL estimators, which we anticipate can be improved.
5 Limitations and Broader Impact
In addition to the limitations that NATURAL shares with every observational study, i.e., the validity of the practitioner’s causal assumptions, it comes with an extra dependence on how well one can approximate the desired conditional distributions. While more and more capable LLMs are being continually developed, the extent to which they satisfy NATURAL’s assumptions is nearly impossible to formally test. Indeed, while pretraining tends to produce calibrated LLM predictions [20], post-training techniques can compromise calibration [31]. Therefore, we emphasize that NATURAL was not developed to recommend therapeutics directly to end-users or to directly inform high-stakes public policies. Instead, we envision NATURAL as a powerful tool to help us approximate ATEs at scale and prioritize confirmatory CREs. We strongly recommend that all predictions made by NATURAL estimators be validated experimentally before being used to inform high-stakes decision-making. Apart from its dependence on LLM capabilities, NATURAL is also limited by the nature of observational, unstructured natural language data:
  • Network Interference. In practice, acquiring i.i.d. reports can be challenging. For instance, social network users might talk to each other and influence their treatment choices. This is a well-known issue in causal inference and statistical sciences in general. Existing solutions rely on a known network structure to sample individuals or correct for their neighbors’ treatments [9, 24, 15].
  • Outcome Measurement. Since NATURAL deals with self-reports, subjects need to be able to report the outcomes of interest. For example, this cannot be applied if the outcome is measured with an expensive, inaccessible test. Therefore, the study design implemented with NATURAL must account for the accessibility of endpoints to users.
  • Selection Bias. Results might be biased towards individuals’ choice of reporting an outcome given their experience with the treatment. Luckily, outcome missingness is a widely studied problem in causality research, see e.g., how to test [8] or how to mitigate [28] it. Note, however, that solutions will often accumulate assumptions on top of NATURAL and should always be critically evaluated by practitioners. Finally, apart from individuals’ choice of reporting, selection bias might also arise from which individuals participate in online forums, i.e., our framework is only capable of estimating local ATEs —external validity is not guaranteed a priori.
6 Empirical Evaluation
Figure 3: For Hillstrom (left) and Retail Hero (right), the KL divergence between estimated joint and propensity distributions and their true counterparts reduces with increasing number of posts (top), as does the RMSE between the NATURAL Full estimate and true ATE (bottom).
Evaluating an end-to-end pipeline for causal inference from unstructured real-world text data to ATEs presents challenges regarding access to data, ground truth ATEs and insightful intermediate metrics. We used two synthetic datasets where we augmented randomized data to mimic real-world observations, while continuing to have access to ground truth evaluation. In addition, we study four real datasets, curated from publicly available Reddit posts from the PushShift dataset, as described in section 4. These six datasets allowed us to systematically evaluate NATURAL.
Table 1: The NATURAL IPW ATE outperforms other versions of the method as well as trained baselines on synthetic datasets, as measured by RMSE.
Hillstrom Retail Hero
[…]
Uncorrected $1.86\pm 0.67$ $4.28$ $0.26\pm 0.30$ $3.08$
N-Full $4.26\pm 0.86$ $2.02$ $1.86\pm 1.38$ $2.08$
N-MC OI $6.17\pm 1.61$ $1.61$ $4.94\pm 2.17$ $2.70$
N-MC IPW $4.81\pm 0.80$ $1.51$ $1.85\pm 2.01$ $2.49$
N-OI $4.58\pm 0.61$ $1.62$ $2.99\pm 1.43$ $1.72$
N-IPW $\mathbf{5.23\pm 1.00}$ $\mathbf{1.32}$ $\mathbf{3.83\pm 1.29}$ $\mathbf{1.39}$
Bag-of-Words $7.57\pm 1.37$ $2.23$ $2.61\pm 2.08$ $2.42$
Sentence Encoder $00.00\pm 0$ 0.00 $6.09$ $1.97\pm 1.62$ $2.10$
IPW (Structured) $6.38\pm 0.26$ $0.39$ $3.09\pm 0.19$ $0.30$
Ground Truth $\mathbf{6.09}$ [17] - $\mathbf{3.32}$ [46] -
Synthetic Datasets.
Causal effect estimation is typically evaluated using synthetic datasets with one or more relationships between the observed covariates, treatment and outcome being contrived. We instead synthesized unstructured observational text data from real randomized tabular datasets, using an LLM. Specifically, we (i) introduced confounding bias by sampling datapoints according to an artificial propensity score, (ii) randomly dropped covariates, (iii) described covariates, treatment and outcome in shuffled orderings, (iv) simulated realism by sampling a persona from the the Big Five personality traits [25] for each datapoint and finally, (v) prompted the LLM to generate a realistic report describing the provided information in the style of someone with the given traits (see appendix D for the full prompt). We used two standard, publicly available randomized datasets: Hillstrom [17] and Retail Hero [46], and plan to open-source scripts to generate our data. Step (i) above is in a similar vein as Keith et al., [22], in that our subsampling strategy does not modify the marginal distribution over covariates and the ATE remains identifiable from observational data.
Real-world Datasets.
To study how our framework may be deployed to test hypotheses using real data from online forums; we considered two medical conditions for which there exist abundant Reddit posts in the PushShift collection [6], with individuals’ personal experiences: the effect of diabetes medications (e.g. Semaglutide) on weight loss and the tolerability of migraine treatments. For each condition, we picked two clinical trials which performed head-to-head comparisons of two treatments that we expected to find references to in relevant subreddits. Moreover, to mitigate selection bias we selected pairs of similar treatments, e.g., comparable availability, where we believe the probability of a user reporting their experience is approximately equal in both. As we will discuss, our results suggest external validity as well, meaning that the probability of a user reporting their experience with the treatments seems to be (approximately) equal to the prior probability of a user reporting any experience. We limited our data collection to posts that were written before December 2022 and made publicly available in the PushShift archives. We curated four datasets for comparison between different treatments, each of which has a ground truth RCT: Semaglutide vs. Tirzepatide [16] and Semaglutide vs. Liraglutide [7] for their effect on weight loss and Erenumab vs. Topiramate [34] and OnabotulinumtoxinA vs. Topiramate [36] for their tolerability. We used the first of these to validate implementation choices NATURAL (like filtering, imputations, prompt specifications) and the other three as held-out test settings, see appendix C. We include further details for all our datasets in appendix E, including the definitions of covariates and outcomes.
Next, we investigate several questions about the performance of NATURAL empirically. We used GPT-4 Turbo for sampling and LLAMA2-70B for computing conditional probabilities.
How well does NATURAL estimate observational distributions from self-reported data?
Our synthetic datasets give us access to the true joint distributions $P(X,T,Y)$ and true propensity scores $P(T=1|X)$. The top row of fig. 3 shows the KL divergence between these distributions and those estimated by NATURAL Full, for Hillstrom (left) and Retail Hero (right). We find that these KL divergences decrease steadily as the number of reports used in the estimation increases. The bottom row shows corresponding root-mean-squared error (RMSE) between NATURAL and the true ATE. This corroborates the insight that as the joint distribution and propensity scores are estimated more accurately, the predicted ATE gets closer to its true value. In particular, we observe a clear correlation between the quality of estimated propensity scores and estimated ATEs.
Tuned Held-out
Semaglutide vs. Tirzepatide (weight loss $\geq$ 5%) Semaglutide vs. Liraglutide (weight loss $\geq$ 10%) Erenumab vs. Topiramate (discontinued due to AE) OnabotulinumtoxinA vs. Topiramate (discontinued due to AE)
[…]
Uncorrected $-33.56\pm 0.77$ $43.67$ $-83.57\pm 0.43$ $68.87$ $29.07\pm 0.48$ $2.87$ $21.55\pm 1.22$ $19.49$
N-MC OI $5.43\pm 1.01$ $4.79$ $-7.71\pm 0.91$ $7.05$ $23.91\pm 1.63$ $4.68$ $46.21\pm 1.94$ $5.55$
N-MC IPW $5.23\pm 0.93$ $4.97$ $-7.43\pm 0.93$ $7.33$ $25.29\pm 1.72$ $3.47$ $46.23\pm 1.93$ $5.57$
N-OI $4.36\pm 2.05$ $6.09$ $\mathbf{-15.90\pm 1.14}$ $\mathbf{1.65}$ $31.21\pm 1.68$ $3.36$ $44.91\pm 1.46$ $4.17$
N-IPW $\mathbf{8.83\pm 0.36}$ $\mathbf{1.33}$ $-12.21\pm 1.09$ $2.72$ $\mathbf{27.90\pm 0.99}$ $\mathbf{1.06}$ $\mathbf{42.60\pm 2.02}$ $\mathbf{2.58}$
Ground Truth $\mathbf{10.11}$ [NCT03987919, 16] $\mathbf{-14.7}$ [NCT03191396, 7] $\mathbf{28.3}$ [NCT03828539, 34] $\mathbf{41.00}$ [NCT02191579, 36]
Table 2: Using real data, best performing NATURAL estimators fall within 3 percentage points of their corresponding ground truth clinical trial ATEs. Possible ATE values lie between $-100$ and $100$.
How do NATURAL methods compare to one another and to trained baselines?
We present our estimated ATE and its RMSE on the synthetic datasets in table 1. Further, we evaluate two trained baselines, which use a Bag-of-Words model and a sentence encoder respectively, to train representations of text data with their labels. Here, for each attribute in the set of covariates, treatments, and outcomes, we train a MLP model with 5-fold cross validation to predict that attribute. We then use these predicted attributes as a tabular dataset of samples that can be plugged into any causal inference estimator. We find that our methods are competitive with or outperform these baselines, despite not being trained with any labels. In particular, the sentence encoder baseline collapsed to an ATE of zero for Hillstrom, having learned the constant predictor for outcomes.
Table 2 compares NATURAL methods to estimate the ATEs in real-world clinical settings using self-reported data from the PushShift collection of Reddit posts. Remarkably, our predicted ATEs (a) depict the same direction of effect, and (b) fall within 3 percentage points of their corresponding ground truth clinical trial ATEs. For both synthetic and real data experiments, NATURAL IPW outperforms other versions across datasets, except for the Semaglutide vs. Liraglutide setting, where NATURAL OI performed the best. Both N-MC versions perform similarly on all datasets.
This result is significant. Clinical trials can take on the order of years and costs in the tens to hundreds of millions of dollars. Going from the raw language observational data to ATE in our framework takes on the order of days and costs at most a few hundred dollars of compute. For problems in medicine, economics, sociology, and political science where randomization is infeasible or expensive, NATURAL provides a tractable way to leverage observational data to rank potential experiments prior to conducting them.
Figure 4: Ablation study on Semaglutide vs. Tirzepatide, to tease apart the effect of data filtering and imputation (left) as well as LLM scale for conditionals (right) on NATURAL performance.
How do different choices in the NATURAL pipeline effect ATE prediction?
We assess the impact of key choices in our pipeline described in section 4, by ablating them one-by-one. We investigated and selected these choices on the Semaglutide vs. Tirzepatide experiment. Figure 4 (left) compares the RMSE of predicted ATEs when data is not filtered according to inclusion criteria and LLM imputations are replaced with samples from an uniform distribution. It shows that both inclusion-based filtering and imputations from a pretrained LLM are crucial for the performance of NATURAL. We also compared performance of our method when the conditional probabilities in eq. 6 are evaluated using models of different scales in fig. 4 (right), and found that performance improves at larger scales and with greater quantity of data.
Figure 5: NATURAL propensity scores balance the Semaglutide vs. Tirzepatide covariates better than uniform scores.
How well do different estimates of propensity score balance covariates?
A property of accurate propensity score estimates is that they balance covariates across treatment cohorts (see Ding, [10] for details and proofs), i.e. the average treatment effect on each covariate, corrected using propensity scores, is close to zero. Figure 5 visualizes this quantity for different covariates of the Semaglutide vs. Tirzepatide experiment and shows that propensity scores estimated using LLAMA2 conditional distributions balance the covariates far better than a uniform distribution does, with the 70B model consistently estimating the treatment effect on each covariate as close to zero. Similar visualizations for the test settings are shown in fig. 10 of section F.2.
7 Conclusion
In this work, we introduced NATURAL, a family of text-conditioned estimators, to automate treatment effect estimation using free-form text data. We demonstrated NATURAL’s efficacy with six synthetic and real datasets for systematic evaluation of its pipeline. We exposed the ability of LLMs to extract meaningful conditional distributions over structured variables and how their combination with classical causal estimators can predict real-world causal effects with remarkable accuracy. Given this promising performance, exciting directions for future work include (i) incorporating automatic prompt tuning methods into the pipeline, (ii) extending our methods to real-valued treatments, outcomes and covariates, (iii) exploring whether our assumptions can be weakened, (iv) exploring other domains in applied research, e.g., social sciences, (v) performing a more extensive evaluation of NATURAL on different study designs to better understand what type of treatments, outcomes, and reports show better or worse practical performance with NATURAL or (vi) deploying the pipeline to test hypotheses at even larger scales.
NATURAL estimators have numerous use cases with potentially far-reaching impact. As long as patients have access to treatments and report their experiences, NATURAL can be used to compare two treatments in new indications or new populations. Therefore, our pipeline can in principle support efforts to prioritize trials for repurposed drugs or supplements in under-served diseases or populations. Further, a crucial step after drug approval is post-marketing surveillance for side effects (positive or negative) that may not have been measured or may have been too rare to identify in a smaller trial. NATURAL can leverage the diversity of available language data to detect these effects. While our motivations largely stem from the challenges of drug development, our NATURAL estimators are applicable to any effect estimation setting for which there exists relevant natural language data.
Acknowledgements
[…]
Appendix
Appendix A Notation
$R$ Random variable corresponding to unstructured natural language text from a social media post (or report).
$X$ Random variable corresponding to features of an individual in a causal inference dataset.
$T$ Random variable corresponding to treatment or intervention assigned to an individual in a causal inference dataset.
$Y$ Random variable corresponding to outcome observed for an individual in a causal inference dataset.
$x$ Possible instance of $X$ from its support $\mathcal{X}$.
$t$ Possible instance of $T$ from its support $\mathcal{T}=\{0,1\}$ (binary treatments).
$y$ Possible instance of $Y$ from its support $\mathcal{Y}=\{0,1\}$ (binary outcomes).
$r$ Possible instance of $R$ from its support $\mathcal{R}$.
$Y(t)$ Random variable corresponding to potential outcome observed for an individual after receiving treatment $t$.
$e(X)$ Propensity score function for binary treatments, equal to $P(T=1|X)$.
$X_{i}$ Sampled value of $X$ for individual $i$.
$T_{i}$ Sampled value of $T$ for individual $i$.
$Y_{i}$ Sampled value of $Y$ for individual $i$.
$R_{i}$ Sampled report $R$ for individual $i$.
$\tau$ Average treatment effect (ATE) given by $\mathbb{E}[Y(1)-Y(0)]$, where the expectation is over some defined population of individuals.
$n$ Total number of individuals.
$n_{1}$ Total number of individuals that are assigned treatment $T=1$.
$n_{0}$ Total number of individuals that are assigned treatment $T=0$.
Appendix B Deriving NATURAL Full from Outcome Imputation estimator
In section 3, we derived the idealized version of our method, NATURAL Full using the form of the IPW estimator. In fact, any unbiased estimator would lead to same form as eq. 3, which can then be used to derive NATURAL Full using reports $R$ and teh law of iterated expectations. Here, we show this using the Outcome Impuation estimator. For discrete $X$ and binary $T$ and $Y$, we have
$\tau$ $=\mathbb{E}_{X}[P(Y=1\mid T=1,X)-P(Y=1\mid T=0,X)]$ (outcome imputation) (11)
$=\sum_{x\in\mathcal{X}}P(X=x)[P(Y=1\mid T=1,X=x)-P(Y=1\mid T=0,X=x)]$ (expectation of discrete $X$) (12)
$=\sum_{x\in\mathcal{X}}P(X=x)\left[\frac{P(Y=1,T=1,X=x)}{P(X=x)P(T=1|X=x)}-\frac{P(Y=1,T=0,X=x)}{P(X=x)P(T=0|X=x)}\right]$ (expanding conditionals) (13)
$=\sum_{x\in\mathcal{X}}\left[\frac{P(Y=1,T=1,X=x)}{e(x)}-\frac{P(Y=1,T=0,X=x)}{1-e(x)}\right]$ (definition of $e(x)$) (14)
$=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}y\left[\frac{P(Y=y,T=1,X=x)}{e(x)}-\frac{P(Y=y,T=0,X=x)}{1-e(x)}\right]$ (since $\mathcal{Y}=\{0,1\}$) (15)
$=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}y\sum_{t\in\mathcal{T}}\left[t\frac{P(Y=y,T=t,X=x)}{e(x)}-(1-t)\frac{P(Y=y,T=t,X=x)}{1-e(x)}\right]$ (since $\mathcal{T}=\{0,1\}$) (16)
$=\sum_{(x,t,y)}P(Y=y,T=t,X=x)\left[\frac{ty}{e(x)}-\frac{(1-t)y}{1-e(x)}\right]$ (rearranging terms) (17)
$=\mathbb{E}_{(X,T,Y)}\left[\frac{TY}{e(X)}-\frac{(1-T)Y}{1-e(X)}\right],$ (18)
which is equivalent to eq. 3.
Appendix C Worked Example: Semaglutide vs. Tirzepatide
To make the NATURAL pipeline and its implementation more concrete, we now work through an end-to-end example using the Semaglutide vs. Tirzepatide dataset. We used this setting to develop our evaluation setup and constructed the pipeline as a function of the experiment design. While it is infeasible to exhaustively describe the entire decision space, major decisions that were found to impact ATE prediction are (i) filtering strategy, (ii) prompt tuning for extraction, and (iii) covariates’ discretization. The strategies we tried are included in the pipeline description below, with the specific decisions made for each highlighted in this color.
Pipeline for Semaglutide vs. Tirzepatide.
Given the experimental design of the clinical trial in Frías et al., [16] (NCT03987919), we defined our experiment as follows:
  1. 1.
    Treatments: Semaglutide and Tirzepatide
  2. 2.
    Outcome: Percentage of participants who lost 5% or more of their initial weight
  3. 3.
    Covariates: Age, Sex, BMI, Start weight, Start HbA1c, Duration
  4. 4.
    Inclusion criteria:
    1. (a)
      The user must be diagnosed with Type 2 Diabetes with starting HbA1c between 7% and 10.5%.
    2. (b)
      They must already be on a regime of the treatment called Metformin.
    3. (c)
      They must have a BMI of 25 kg/m2 or more.
    4. (d)
      Since different treatment dosages can have varying effects, we also included dosage as an inclusion criterion, i.e. we aimed to include only posts that reported taking 1mg for Semaglutide and 5mg for Tirzepatide, as in the clinical trial.
Decisions in the pipeline that are a direct function of an experimental design like the one above are highlighted in this color in the following description.
Next, we implemented the entire pipeline as follows:
  1. 1.
    Initial filter. We identified nine subreddits relevant to this problem setting: r/Mounjaro, r/Ozempic, r/fasting, r/intermittentfasting, r/keto, r/loseit, r/Semaglutide, r/SuperMorbidlyObese, r/PlusSize. From each subreddit, we downloaded all submissions and comments posted upto December 2022 from the PushShift collection, so as to only use publicly available data. This resulted in a dataset of 577,733 submissions and comments. An initial deterministic, task-agnostic and rule-based filter removed any submission or comment if its content was not a string, if it had no score, if the content was "[deleted]" or "[removed]", if it was a comment with fewer than ten space-separated strings (presumably, words), if the author’s name contained the string "bot", if there were no spaces in the first 2048 characters, and if less than 50% of all characters were alphabetic. This reduced the dataset size to 380,276. We then formatted this data into dictionary-like datapoints with fields: subreddit, title, date created, post/comment, author replies. We indcluded the last field because comments written by the author as replies to their own post may contain additional relevant information when combined with with original post and other replies. We then passed these through a task-dependent string-matching filters. For this dataset in particular, we listed strings used commonly to refer to the treatments, ["ozempic", "mounjaro", "semaglutide", "tirzepatide", "wegovy", "rybelsus", "zepbound"], included common misspellings generated with GPT-4 and Perplexity, and filtered out datapoints that did not contain any of these strings. Similarly, we listed keywords relevant to the outcome of interest, ["kg", "kilo", "lb", "pound", "weigh", "drop", "loss", "lost", "gain", "hb", "a1c", "hemoglobin", "haemoglobin", "glucose", "sugar"] and filtered out datapoints that did not contain any of these strings. This filtered dataset now contained 50,654 datapoints.
  2. 2.
    Filter by relevance. Next, we wrote a problem setting description and prompted GPT-3.5 Turbo to determine whether the posts, along with auxiliary information from the formatted dictionaries described above, were relevant to the described setting. The description and instructions for this particular dataset are shown in appendix D. We manually labeled a handful of datapoints as Yes or No and included these as incontext examples to improve the LLM’s generations. We removed datapoints that were deemed irrelevant, resulting in a "relevant" dataset of 21,229 datapoints.
  3. 3.
    Filter by treatment-outcome. To further filter the data to points that refer specifically to the treatments and outcome of interest, we prompted GPT-3.5-Turbo to extract only information required to ascertain the treatment and outcome, as shown in appendix D. Since the outcome for this dataset, achievement of a target weight loss of 5% or more, may be reported in several ways, we attempted to cover all those possibilities. Specifically, we prompted the LLM to extract the user’s starting weight, end weight, change in weight and percentage of change in weight. Several combinations of these attributes allow us to programmatically infer the final outcome. We also extracted the units in which weight was reported, converting all extractions to be in lbs. We filtered out any datapoint for which the extracted treatment was not one of the treatments considered for this task or for which it was not possible to infer the outcome using the above-mentioned extracted information. This finally gave us a natural language dataset of 4619 relevant reports, each of which contained treatment and outcome information pertaining to the defined problem setting.
  4. 4.
    Filter known covariates by inclusion criteria. To evaluate against the clinical trial, we further filtered the dataset according to its inclusion criteria. At a high level, different strategies for this filtering trade-off how strictly we match the criteria with how many datapoints remain in the filtered set. Strict filtering to match every criteria exactly resulted in very few reports. Less strict filtering to match a subset of the criteria resulted in predicted ATEs that varied depending on which criteria we chose to satisfy, with no apparent task-agnostic principle to determine crucial criteria. Finally, we aimed for a set of reports with non-zero probability of satisfying these criteria. As described in item (iv) of section 4 and motivated in appendix G, we extracted all covariates including ones related to the inclusion criteria and removed datapoints whose extractions were not Unknown and failed to satisfy the criteria above. This resulted in a dataset of 1265 reports.
  5. 5.
    Extract known and unknown covariates. Treating these 1265 reports as the final dataset from which to estimate an ATE, we extracted the set of covariates given in our experiment definition. We also included the duration of treatment as a covariate since this information is often reported and is likely to influence the outcome. This extraction step was conditioned on inclusion criteria being satisfied, a description of which was included in the extraction prompt, as in appendix D. We tuned the prompts for extracting attributes, which include general instructions for the task and specific questions for each attribute. This was done by inspecting a handful of reports and corresponding extractions and then modifying prompts to correct any observed errors.
  6. 6.
    Infer conditionals. We inferred conditional distributions from LLAMA2-70B for different versions of NATURAL, with the strategy described in item (vi) of section 4 and LLM inputs of the form shown in appendix D. Here, "conditioning on covariates" was implemented by adding questions about the covariates and their sampled answers to the input. For instance, for sex, the question "What is the reported sex of the user?" was followed by its previously extracted answer (Male or Female). The scoring strategy required enumerating possible options for treatments and outcomes for each input, which were ["Semaglutide like Ozempic or Wegovy or Rybelsus", "Tirzepatide like Mounjaro or Zepbound"] and ["No", "Yes"], respectively.
  7. 7.
    Finally, given all the required extractions and conditional probabilities, we required discrete covariates to plug them into our NATURAL estimators. Hence, we converted any continuous covariates into discrete categories. These categories for each dataset are shown in table 4 for all our datasets. Different choices of discretization led to slightly different ATE predictions. We found it most helpful to discretize continuous numerical covariates into intervals such that the number of datapoints were roughly balanced across intervals. This avoided covariate strata with too many or too few datapoints and resulted in ATE predictions from all NATURAL estimators that were sufficiently close to the ground truth.
Adapting the pipeline to test trials.
The decisions above in this color directly or indirectly influenced the ATE and we made our choices with access to the ground truth for Semaglutide vs. Tirzepatide. Hence, we call this a "tuned" setting. We fixed these decisions for the three test settings. Note that no other aspect of the pipeline depends on the ATE. All the choices in this color are a function of the experimental design. Hence, the pipeline can be easily adapted to any new setting, given its experimental design and without knowledge of the true ATE.
Appendix D LLM Prompts
Prompt 1: Synthetic report generation (Hillstrom) Prompt 2: Relevance filtering (Weight Loss) Prompt 3: Treatment-outcome filtering (Weight Loss) Prompt 4: Covariate extraction (Weight Loss) Prompt 5: Covariate imputation (Weight Loss) Prompt 6: Conditional distribution inference (Weight Loss)
Appendix E Dataset Details
We provide further details about the treatments, outcomes and covariates, along with inclusion criteria and discrete categories used in our experiments, for each dataset in tables 3 and 4.
Dataset Treatment Outcome Synthetic confounder
Hillstrom email communication website visit newbie
Retail Hero SMS communication purchase age
Semaglutide vs. Tirzepatide corresponding drug weight loss of 5% or more NA
Semaglutide vs. Liraglutide corresponding drug weight loss of 10% or more NA
Erenumab vs. Topiramate corresponding drug discontinuation due to adverse effects NA
OnabotulinumtoxinA vs. Topiramate corresponding drug discontinuation due to adverse effects NA
Table 3: Treatments, outcomes and synthetic confounders (where applicable) for each dataset.
Covariate Description Discrete categories Inclusion criteria
Hillstrom
recency number of months since last purchase [$1-4,5-8,9-12$] NA
history dollar value of previous purchase [$0$ 100 , 100 - 200 , … , > 1000]
mens purchase of men’s merchandise [True,False]
womens purchase of women’s merchandise [True,False]
zip_code type of area of residence [Suburban area,Rural area,Urban area]
newbie new customer [True,False]
channel channel used for purchases [Phone,Web,Multichannel]
Retail Hero
avg. purchase avg. purchase value per transaction [1263,264396,397611,>6121-263,264-396,397-611,>6121 - 263 , 264 - 396 , 397 - 611 , > 612] NA
avg. product quantity avg. number of products bought [7,>7$\leq$ 7,>7≤ 7 , > 7]
avg. points received avg. number of points received [5,>5$\leq$ 5,>5≤ 5 , > 5]
num transactions total number of transactions so far [8,915,1627,>28$\leq$ 8,9-15,16-27,>28≤ 8 , 9 - 15 , 16 - 27 , > 28]
age age of user [45,>45$\leq$ 45,>45≤ 45 , > 45]
Semaglutide vs. Tirzepatide
age age of user [45,>45$\leq$ 45,>45≤ 45 , > 45] (t2 diabetes==True) & (7 $\leq$ start HbA1c $\leq$ 10.5)) & (metformin==True) & (bmi $\geq$ 25)
sex sex of user [Male,Female]
bmi body mass index of user [28.5,>28.5$\leq$ 28.5,>28.5≤ 28.5 , > 28.5]
start HbA1c initial glycated haemoglobin value [7.5,>7.5$\leq$ 7.5,>7.5≤ 7.5 , > 7.5]
start weight initial weight in lbs [220,>220$\leq$ 220,>220≤ 220 , > 220]
duration (days) number of days treatment was taken for [90,>90$\leq$ 90,>90≤ 90 , > 90]
Semaglutide vs. Liraglutide
age age of user [45,>45$\leq$ 45,>45≤ 45 , > 45] (t2 diabetes==True) & (7 $\leq$ start HbA1c $\leq$ 11)) & (metformin/other==True)
sex sex of user [Male,Female]
bmi body mass index of user [28.5,>28.5$\leq$ 28.5,>28.5≤ 28.5 , > 28.5]
start HbA1c initial glycated haemoglobin value [7.5,>7$\leq$ 7.5,>7≤ 7.5 , > 7]
start weight initial weight in lbs [220,>220$\leq$ 220,>220≤ 220 , > 220]
duration (days) number of days treatment was taken for [120,>120$\leq$ 120,>120≤ 120 , > 120]
Erenumab vs. Topiramate
age age of user [32,>32$\leq$ 32,>32≤ 32 , > 32] (18 $\leq$ age $\leq$ 65) & (pregnant==False) & (baseline MMD $\geq$ 4)
sex sex of user [Male,Female]
country country of residence [United States,Canada,…]
baseline MMD initial number of monthly migraine days [6,>6$\leq$ 6,>6≤ 6 , > 6]
duration (days) number of days treatment was taken for [30,>30$\leq$ 30,>30≤ 30 , > 30]
OnabotulinumtoxinA vs. Topiramate
age age of user [25,>25$\leq$ 25,>25≤ 25 , > 25] (18 $\leq$ age $\leq$ 65) & (baseline MMD $\geq$ 15)
sex sex of user [Male,Female]
country country of residence [United States,Canada,…]
baseline MMD initial number of monthly migraine days [15,>15$\leq$ 15,>15≤ 15 , > 15]
duration (days) number of days treatment was taken for [30,>30$\leq$ 30,>30≤ 30 , > 30]
Table 4: Covariate descriptions, corresponding discrete categories and inclusion criteria enforced for each dataset. Intervals for continuous numerical variables were determined from the extracted values such that each discrete category is roughly balanced in terms of its number of datapoints.
Appendix F Further Experimental Results
F.1 Known and Unknown/Imputed covariates for real data experiments
We refer the reader to figs. 6, 7, 8 and 9 for empirical distributions of covariates extracted by an LLM in its first extraction as well as those imputed in its second imputataion conditioned on inclusion criteria.
Figure 6: Distributions of "known" (top) vs "unknown" and imputed (bottom) covariates for Semaglutide vs. Tirzepatide.
Figure 7: Distributions of "known" (top) vs "unknown" and imputed (bottom) covariates for Semaglutide vs. Liraglutide.
Figure 8: Distributions of "known" (top) vs "unknown" and imputed (bottom) covariates for Erenumab vs. Topiramate.
Figure 9: Distributions of "known" (top) vs "unknown" and imputed (bottom) covariates for OnabotulinumtoxinA vs. Topiramate.
F.2 Balancing property of propensity scores
We refer the reader to Figure 10 for visualizations of the propensity score corrected average treatment effect on covariates for all test clinical settings. For each setting, our estimated propensity score balances each covariate, far better than a uniform propensity distribution would.
Since covariates may take values at different scales, we computed the standard mean difference (SMD) across cohorts for each covariate $X^{(i)}$ [14], given by:
$SMD=\frac{X^{(i)}(1)-X^{(i)}(0)}{\sqrt{0.5*(\texttt{var}(X^{(i)}(1))+\texttt{var}(X^{(i)}(0)))}},$ (19)
where $X^{(i)}(1)-X^{(i)}(0)$ estimates the average treatment effect on $X^{(i)}$, using propensity score weighting, and $\texttt{var}(\cdot)$ denotes sample variance.
(a) Semaglutide vs. Liraglutide
(b) Erenumab vs. Topiramate
(c) OnabotulinumtoxinA vs. Topiramate
Figure 10: Propensity scores estimated with LLAMA2-70B balance covariates of the real clinical datasets, far better than uniform scores do.
Appendix G Inclusion Criteria conditioned Estimator
We are interested in an ATE conditioned on inclusion criteria denoted $I$,
$\tau(I)=\mathbb{E}[Y(1)-Y(0)\mid X\in I].$ (20)
Let $\tau(X,T,Y)$ be a function such that
$\tau(I)=\mathbb{E}_{X,T,Y}[\tau(X,T,Y)\mid X\in I].$ (21)
For example, if $\tau(I)$ can estimated by the IPW estimator,
$\tau(X,T,Y)=\frac{TY}{e(X)}-\frac{(1-T)Y}{1-e(X)},$
because the $P(T=1|X=x,X\in I)=P(T=1|X=x)$. Throughout this section, we operate under Assumptions 3 and 4 and assume that the LLM gives us access to the true data-generating conditionals.
The law of total expectation gives us an estimator that can operate on samples of reports $R$:
$\tau(I)$ $=\mathbb{E}_{X,T,Y}[\tau(X,T,Y)\mid X\in I]$
$=\mathbb{E}_{R|X\in I}[\mathbb{E}_{X,T,Y}[\tau(X,T,Y)\mid X\in I,R]]$
$=\sum_{r}P(R=r|X\in I)\mathbb{E}_{X,T,Y}[\tau(X,T,Y)\mid X\in I,R]$
$=\sum_{r}P(R=r)\frac{P(X\in I|R=r)}{P(X\in I)}\mathbb{E}_{X,T,Y}[\tau(X,T,Y)\mid X\in I,R]$
$=\mathbb{E}_{R}\left[\frac{P(X\in I|R)}{P(X\in I)}\mathbb{E}_{X,T,Y}[\tau(X,T,Y)\mid X\in I,R]\right].$
To summarize, we have the identities:
$\tau(I)$ $=\mathbb{E}_{R|X\in I}[\mathbb{E}_{X,T,Y}[\tau(X,T,Y)\mid X\in I,R]]$ (22)
$=\mathbb{E}_{R}\left[\frac{P(X\in I|R)}{P(X\in I)}\mathbb{E}_{X,T,Y}[\tau(X,T,Y)\mid X\in I,R]\right].$ (23)
It is possible to directly estimate eq. 23 with Monte Carlo samples from $P(R)$ and an approximation of $P(X\in I|R)/P(X\in I)$ using an LLM, but the latter is computationally very expensive. Alternatively, one can use rejection sampling to simulate $R|X\in I$ by sampling $X_{i}\sim P(X|R_{i})$ and rejecting $R_{i}$ if $X_{i}\notin I$. This step could be very wasteful for precious data. In fact, in our experiments with the Semaglutide vs. Tirzepatide dataset, we found that using GPT-4 Turbo for sampling $X_{i}\sim P(X|R_{i})$ resulted in a very peaky distribution with little diversity in the extractions, even after increasing its temperature argument as much as possible without sacrificing generation quality. As a result, after rejecting $R_{i}$ if $X_{i}\notin I$, we were left with a very small number of reports, that would be infeasible to plug into any estimator. Therefore, we devise a few assumptions that allow us to estimate eq. 21 without these approaches.
Let $X\in\mathbb{R}^{D}$ and let us make the following assumption on the inclusion criteria:
Assumption 5 (Inclusion criteria specification)
The inclusion criterion $I$ defines a box, i.e., it is specified separately for each covariate dimension $I^{d},d\in\{1,\dots,D\}$ and the set of covariates satisfying every inclusion criteria is given by the product of individual criteria over the covariate dimensions, i.e., $\{X\in I\}=\prod_{d=1}^{D}\{X^{d}\in I^{d}\}$ where $X^{d}$ is the $d$-th dimension of $X=(X^{d})_{d=1}^{D}$.
Recall from section 4 that inclusion-based filtering leaves us with reports whose covariates are either “known” and satisfy their criteria or Unknown. We also have the value of the known covariates. Let $K\in\{0,1\}^{D}$ be the binary vector of variables $K^{d}$ that indicate whether the covariate $X^{d}$ is found to be “known” for a random report $R$. Let $X^{K}=(X^{d}:K^{d}=1)$ be the vector of length $\sum_{d}K^{d}$ holding the values of the known covariates. For ease of notation, define the event that the known covariates satisfy their criteria and the event that the unknown covariates satisfy their criteria:
$\{X^{K}\in I^{K}\}$ =$=\{X^{d}\in I^{d},\forall d\colon K^{d}=1\}$ (24)
$\{X^{1-K}\in I^{1-K}\}$ =$=\{X^{d}\in I^{d},\forall d\colon K^{d}=0\}$ (25)
Notice that $\{X^{K}\in I^{K}\}\cap\{X^{1-K}\in I^{1-K}\}=\{X\in I\}$. Thus, after the filtering steps we have
$\{R_{i},K_{i},X_{i}^{K_{i}}\}_{i=1}^{n}$ (26)
with the guarantee that the knowns satisfy their inclusion criteria, $\{X^{K_{i}}_{i}\in I^{K_{i}}\}$. Formally, assuming that the LLM computes the true conditional distribution of the data-generating process (Assumption 4), this gives us data sampled i.i.d. from $P(R=r,K=k,X^{k}=x^{k}|X^{K}\in I^{K})$. Note, that we are assuming the existence of an additional ground-truth random variable $K$ in the data-generating process that describes whether a covariate is knowable from a report. Here, we show how to estimate $\tau(I)$ from this dataset of filtered reports using importance sampling, under the following assumption:
Assumption 6 (Satisfaction of $I$ by Unknown covariates)
Satisfaction of inclusion criteria by unknown covariates is conditionally independent of the report and the known covariates given satisfaction of inclusion criteria by known covariates, i.e., for all $r,k,x^{k}$:
$P(R=r,K=k,X^{k}=x^{k}|X^{K}\in I^{K},X^{1-K}\in I^{1-K})=P(R=r,K=k,X^{k}=x^{k}|X^{K}\in I^{K})$ (27)
One can derive the following identity in a similar fashion as eq. 23
$\tau(I)=\mathbb{E}_{R,K,X^{K}|X^{K}\in I^{K}}\left[\frac{P(R,K,X^{K}|X\in I)}{P(R,K,X^{K}|X^{K}\in I^{K})}\mathbb{E}_{X,T,Y}[\tau(X,T,Y)\mid X\in I,R,K,X^{K}]\right].$ (28)
From assumption 6, the fraction above simplifies to $1$, leaving us with the following estimator
$\tau(I)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{X_{i},T_{i},Y_{i}}[\tau(X_{i},T_{i},Y_{i})\mid X_{i}\in I,R_{i},K_{i},X_{i}^{K_{i}}],$ (29)
which can be computed from the information available at the end of filtering. In practice, we do not condition the LLM on $K_{i}$ in the final inference step (vi), which amounts to an additional conditional independence assumption:
Assumption 7 (Conditional independence of knowable covariates)
$K\perp\!\!\!\perp(T,Y)\mid(X,R)$.
Equation 29 above can now be more efficiently estimated by prompting the LLM to extract covariates under the constraints of the inclusion criteria for each report in our filtered dataset, and then following the remaining steps in the pipeline to an ATE estimate.